OpenStudy (anonymous):

a displacement is related to the time as x=Asin(2[pi]ft), where A and f are constants. find the dimensions of A. (HINT: A trig function appearing in an equation must be dimensionless.)

OpenStudy (unklerhaukus):

\[x[\text{m}]=A[\text{m}]\sin(2\pi\times f[\text{Hz}] \times t[\text{s}])\]

OpenStudy (unklerhaukus):

\[[\text{Hz}] =[\text{s}]^{-1}\]

OpenStudy (anonymous):

is this simple harmonic motion? o.O

OpenStudy (unklerhaukus):

yes